# GOUHFI 2.0: A Next-Generation Toolbox for Brain Segmentation and Cortex Parcellation at Ultra-High Field MRI

### v1

@@ 3.2 SCAIFIELD: Comparing GOUHFI’s variants in healthy controls at 7T
A coronal slice of an N4-MPRAGE image from two healthy subjects in the SCAIFIELD dataset is shown in Figure 5 for all segmentation techniques tested in this work. For subject A, the delineation of cerebellum WM and cortex is shown, with a special focus on the inferior cerebellum WM branches. All GOUHFI variants demonstrated an improved identification compared to FastSurferVINN and SynthSeg+, with a slight but detectable improvement for GOUHFI 2.0-n2. Moreover, GOUHFI and SynthSeg+ over-detection of superior cerebellum WM (orange arrows) appeared to be substantially reduced with both GOUHFI 2.0 variants. For subject B, although the inferior cerebellar white matter branches are still visible (yellow arrows), particular attention is drawn to the delineation of the inferior cerebellar cortex (red arrows). The GOUHFI 2.0-n1 variant was the only technique that repeatedly faced problems to properly delineate the inferior cerebellum cortex. That issue was not present in the original GOUHFI. Overall, GOUHFI 2.0-n2 demonstrated the best cerebellum WM branches identification while not being affected by the suboptimal cerebellum cortex segmentation, as was the case for GOUHFI 2.0-n1.
Figure 5: Figure demonstrating the difference in cerebellum segmentations between FastSurferVINN, SynthSeg+, GOUHFI, GOUHFI 2.0-n1 and GOUHFI 2.0-n2 computed from the N4-corrected MPRAGE image for two healthy subjects in the SCAIFIELD dataset. For subject A, the green horizontal lines correspond to the most inferior cerebellar WM branches detected by GOUHFI 2.0-n2 for both hemispheres. The lines for GOUHFI 2.0-n2 are reproduced for every method for comparison. The orange arrows show errors in cerebellar WM segmentation. For subject B, red arrows show the mislabeling of the cerebellar cortex by GOUHFI 2.0-n1 compared to the other techniques. Yellow arrows compare the segmentation of one cerebellar WM branch between techniques.
In order to quantitatively assess the differences in segmentation between the original GOUHFI and the two new variants, the median and interquartile range (IQR) for the DSC and ASD for several brain regions for the N4-MPRAGE and MPRAGElike [undefaab] were computed, and are listed in Table 2. The two variants of GOUHFI 2.0 showed highly similar performance, markedly outperforming SynthSeg+ for all labels with a difference of 6 DSC points on average and close to double the ASD values. GOUHFI 2.0-n2 has the highest DSC and ASD values compared to GOUHFI.
@@ 3.3 SCAIFIELD-NPC: Assessment of segmentation methods in SCA subjects
While the previous results section used healthy subjects at 7T, the performance of GOUHFI 2.0-n2 against FastSurferVINN and SynthSeg+ for three SCA subtypes are shown in Figure 6 with a special focus on the cerebellum. All techniques struggled to accurately segment the cerebellar WM and cortex for the three SCA subjects. While FastSurferVINN had to best overall delineation of the cortex-CSF boundary, it systemically missed some inferior sections for all subjects. Moreover, inferior WM seemed particularly challenging to identify for FastSurferVINN compared to the two other techniques. Conversely, SynthSeg+ and GOUHFI 2.0-n2 over-segmented the superior cerebellar WM, with SynthSeg+ exhibiting a more pronounced error. Overall, SynthSeg+ successfully identified all cerebellar cortex regions, but consistently over-segmented them across all test cases. For the second subject, GOUHFI 2.0-n2 mislabeled some brain cortex voxels as cerebellum cortex.
@@ 3.4 TIV estimation
The results from the same volumetric analysis as done in [undefaaa] are shown in Figure 8 A). Overall, the same consistent decrease trend between HC and PDP as observed for the original GOUHFI for the putamen, hippocampus and amygdala was observed for GOUHFI 2.0-n2. It was only for putamen that all techniques presented a statistically significant difference between both HC and PDP sub-groups. For SynthSeg+, the median volumes measured were systematically larger than the rest. The p-values calculated are reported in Table 3. In addition, since TIV estimation is part of the GOUHFI 2.0 toolbox, an independent volumetry analysis with normalized volume values were performed without the use of SPM12. The same volumetric analysis, using TIV estimates from GOUHFI2.0-n2 and SynthSeg+, is shown in Figure 8 B), with corresponding p-values reported in Table 3. Albeit smaller than the normalized volumes reported using SPM12, the new normalized volumes resulted in the same statistical significance for the putamen. For both FastSurferVINN and the original GOUHFI, the TIV was not calculated independently due to differing definitions for the CSF label. An additional figure showing the volumetry results for the original GOUHFI, GOUHFI 2.0-n2 with TIV from SPM12 and GOUHFI 2.0-n2 with independent TIVs is provided in Appendix 6.2.
Figure 8: Box plots showing the normalized volumes computed by FastSurferVINN (left), GOUHFI 2.0-n2 (middle) and SynthSeg+ (right) for healthy controls (HC) and Parkinson’s disease patients (PDP) for the putamen, hippocampus and amygdala. In A), SPM12 was used for estimating the TIV for normalizing the volumes. In B), the same results are shown but GOUHFI 2.0-n2 and SynthSeg+ were used to estimate the TIV for each respectively (SPM12 was used for FastSurferVINN). For putamen, the three techniques had a statistically significant difference in volume between HC and PDP after Bonferroni correction for both A) and B).
@@ 4 Discussion
While testing the original GOUHFI [undefaaa], incidental observations of the limitations highlighted in Figure 4 prompted the idea of enhancing the diversity of the training dataset. One known limitation of the original training dataset was the absence of aged and/or non-healthy brain anatomies. The original GOUHFI exhibited limitations not observed for FastSurferVINN or SynthSeg+, likely due to the absence of non-healthy anatomies in its training dataset, which were included in the latter two. In order to address this issue, 32 aged and demented subjects were incorporated based on an ad hoc visual assessment of subjects from the OASIS1 dataset and only selecting subjects exhibiting substantially enlarged ventricles. Ultimately, segmentation errors like the ones observed in the original GOUHFI were completely resolved when qualitatively assessing GOUHFI 2.0’s output segmentations for the same test subjects, while also being absent in all other test subjects assessed.
Additionally, this limited capacity to delineate enlarged ventricles also highlighted a limitation of the generative model, which is its incapacity to generate synthetic images with anatomical features observed in non-healthy brains. Until further developments on the DR approach for brain segmentation are proposed, using an anatomically-varied training label maps is highly recommended. As outlined in [undefaah], recent developments in diffusion and flow-based DL models show high potential for improving the image synthesis models, hopefully contributing to a more robust and varied synthesizing process.
@@ 4.2 SCAIFIELD: Comparing GOUHFI’s variants in healthy controls at 7T
Figure 5 presented the unexpected weakness of GOUHFI 2.0-n1 struggling with proper identification of inferior cerebellar cortex, exclusively observed for that model. Surprisingly, this suboptimal cerebellar cortex segmentation for GOUHFI 2.0-n1 did not seem to arise from typical inhomogeneities observed at UHF, since this was mostly observed in homogeneous N4-corrected pTx images (as in Figure 5) while not being observed in inhomogeneous 1 Tx 7T images. Ultimately, one possible explanation for this unexpected behavior for GOUHFI 2.0-n1 could be the randomness involvement in (1) the generation of synthetic images and (2) the initial seeding of the DL training process as shown in [undefap].
The superior identification of cerebellar WM branches by GOUHFI 2.0-n2 as shown for both subjects in Figure 5 makes it a tool of choice for researchers working with the cerebellum, outperforming FastSurferVINN and SynthSeg+. Improvements are also evident when comparing the original GOUHFI to the latest two variants of GOUHFI 2.0. As discussed in [undefaaa], even though SynthSeg+ also uses a DR approach, the difference in training resolutions (1.0 mm3 vs (0.7 mm)3) seems especially important for detecting small regions such as thin cerebellar WM branches.
@@ 4.3 SCAIFIELD-NPC: Assessment of segmentation methods in SCA subjects
Overall, no segmentation technique demonstrated a clear superiority over the others. It is important to mention that this section was focused on the cerebellum only. First, FastSurferVINN performed relatively well considering that these test cases were outside its optimized domain (i.e., (0.6 mm)3 7T images). It was the highest performing technique to properly delineate regions with substantial cortex atrophy. This superior delineation in atrophied cortical regions was most pronounced in the superior part of the cerebellum where both SynthSeg+ and GOUHFI 2.0-n2 failed at accurately identifying the CSF-cortex border (cf. results for subject 1). For subjects 2 and 3 with less atrophy, this was less pronounced. This result can be surprising considering that both DR approaches were trained using perfectly aligned training image and label maps from the synthetic images produced by the generative model from the label maps. In other words, one could have expected a better identification of the CSF-GM border by SynthSeg+/GOUHFI 2.0 since the two regions have high contrast between each other. In the end, this circles back to the previous section discussing that a variation of anatomies inside the training corpus is still important, even when using synthetic training images. However, for subject 3, GOUHFI 2.0-n2 demonstrated a great capacity at accurately identifying the CSF-cortex border, except for the superior region.
@@
On the other hand, FastSurferVINN considerably underestimated the inferior extent of the cerebellar WM in two of the three test subjects while also systematically failing at properly delineating the inferior cerebellar cortex, possibly due to the reduced signal due to inhomogeneities. Both issues can be problematic in volumetric studies where under-estimations of both WM and cortex volumes due to segmentation errors can bias the results. For SynthSeg+, the users should be aware of the opposite trend: over-estimation of WM and cortex volumes. As discussed in the last paragraph, SynthSeg+ wrongly segmented a substantial amount of cortex and CSF as cerebellum WM. For the cortex, SynthSeg+ performed the worst at identifying small atrophy regions, resulting in the cortex label to be segmented as a oversimplified continuous shape instead of following the ”branchy” nature of the atrophied cerebellum cortex for SCA subjects. For GOUHFI2.0-n2, the marginal WM overestimation and improved cortex–CSF border identification, compared to SynthSeg+, indicate that it is a better alternative for high-resolution UHF images.
@@ 4.4 TIV estimation
One interesting result from the TIV estimations was that the only cases where the GOUHFI 2.0 variants under-estimated the TIV, compared to SPM12, are the two HCP contrasts. In contrast, SynthSeg+ exhibited a systematic overestimation of the TIV for the HCP dataset and, more broadly, across all ten test datasets. One difference compared to all other test datasets was that the brain extraction step was not performed as part of this work, since it was already performed as part of the HCP preprocessed data. In both cases, the brain extraction process was very accurate, with no inclusion of extra tissue. For STRAT-PARK, only brain-extracted images were provided, with Nighres [undefaai] used for the brain extraction step. For most subjects in that dataset, a substantial amount of extra-cerebral tissue was still present in the brain-extracted images as shown in Figure 4. For the remaining datasets, the brain extraction step was performed with the ANTsPyNet brain extraction tool implemented as part of the GOUHFI toolbox. The results were visually assessed, and cases with minor extra-cerebral tissue were still retained (negligible compared to START-PARK), as this had no impact on brain segmentation aside from the CSF label. Indeed, for GOUHFI2.0-n2, visual inspection revealed that the CSF label often incorporated extra-cerebral tissue when brain extraction quality was low. On the other hand, GOUHFI 2.0-n1 was not as affected by extra-cerebral tissue, resulting in less TIV overestimation. While the GOUHFI 2.0-n1 variant resulted in higher-quality TIV estimates than GOUHFI 2.0-n2, the latter was still selected as the superior variant due to the reasons stated in section 4.2.
@@
While FastSurferVINN was not considered the reference for earlier brain segmentation quantitative comparison in this study, it was selected as the reference for the volumetry analysis. Based on extensive visual inspection of the segmentation results, FastSurferVINN was deemed as the technique with the most accurate delineation of the hippocampus-ventricle laterally. However, in subjects with strong signal inhomogeneities, medial portions of both the hippocampus and amygdala may be excluded, causing lower estimated volumes and possibly explains of the lower estimates for in Figure 8 A) compared to GOUHFI 2.0-n2 and SynthSeg+.
@@ 4.7 Limitations
The main limitation of the GOUHFI 2.0 toolbox is the brain extraction requirement. As already discussed in the original publication of GOUHFI, most segmentation techniques do not share this requirement and is partly explained by lack of robust whole-head segmentation techniques for UHF images [undefaaa]. Future advances in whole-head segmentation algorithms for UHF images with low signal outside the head, or the use of generative diffusion models to generate synthetic head labels, could help address this issue.

### v2

@@ 3.2 SCAIFIELD: Comparing GOUHFI’s variants in healthy controls at 7T
A coronal slice of an N4-MPRAGE image from two healthy subjects in the SCAIFIELD dataset is shown in Figure 5 for all segmentation techniques tested in this work. For subject A, the delineation of cerebellum WM and cortex is shown, with a special focus on the posterior cerebellum WM branches.
All GOUHFI variants demonstrated an improved identification compared to FastSurferVINN and SynthSeg+, with a slight but detectable improvement for GOUHFI 2.0-n2. Moreover, GOUHFI and SynthSeg+ over-detection of superior cerebellum WM (orange arrows) appeared to be substantially reduced with both GOUHFI 2.0 variants. For subject B, although the posterior cerebellar white matter branches are still visible (yellow arrows), particular attention is drawn to the delineation of the posterior cerebellar cortex (red arrows). The GOUHFI 2.0-n1 variant was the only technique that repeatedly faced problems to properly delineate the posterior cerebellum cortex. That issue was not present in the original GOUHFI. Overall, GOUHFI 2.0-n2 demonstrated the best cerebellum WM branches identification while not being affected by the suboptimal cerebellum cortex segmentation, as was the case for GOUHFI 2.0-n1.
Figure 5: Figure demonstrating the difference in cerebellum segmentations between FastSurferVINN, SynthSeg+, GOUHFI, GOUHFI 2.0-n1 and GOUHFI 2.0-n2 computed from the N4-corrected MPRAGE image for two healthy subjects in the SCAIFIELD dataset. For subject A, the green horizontal lines correspond to the most inferior posterior cerebellar WM branches detected by GOUHFI 2.0-n2 for both hemispheres. The lines for GOUHFI 2.0-n2 are reproduced for every method for comparison. The orange arrows show errors in cerebellar WM segmentation. For subject B, red arrows show the mislabeling of the cerebellar cortex by GOUHFI 2.0-n1 compared to the other techniques. Yellow arrows compare the segmentation of one cerebellar WM branch between techniques.
In order to quantitatively assess the differences in segmentation between the original GOUHFI and the two new variants, the median and interquartile range (IQR) for the DSC and ASD for several brain regions for the N4-MPRAGE and MPRAGElike [14] were computed, and are listed in Table 2. The two variants of GOUHFI 2.0 showed highly similar performance, markedly outperforming SynthSeg+ for all labels with a difference of 6 DSC points on average and close to double the ASD values. GOUHFI 2.0-n2 has the highest DSC and ASD values compared to GOUHFI.
@@ 3.3 SCAIFIELD-NPC: Assessment of segmentation methods in SCA subjects
While the previous results section used healthy subjects at 7T, the performance of GOUHFI 2.0-n2 against FastSurferVINN and SynthSeg+ for three SCA subtypes are shown in Figure 6 with a special focus on the cerebellum. All techniques struggled to accurately segment the cerebellar WM and cortex for the three SCA subjects. While FastSurferVINN had the best overall delineation of the cortex-CSF boundary, it systematically missed some posterior sections of the cerebellum for all subjects. Moreover, posterior WM seemed particularly challenging to identify for FastSurferVINN compared to the two other techniques. Conversely, SynthSeg+ and GOUHFI 2.0-n2 over-segmented the anterior cerebellar WM, with SynthSeg+ exhibiting a more pronounced error. Overall, SynthSeg+ successfully identified all cerebellar cortex regions, but consistently over-segmented them across all test cases. For the second subject, GOUHFI 2.0-n2 mislabeled some brain cortex voxels as cerebellum cortex.
@@ 3.4 TIV estimation
The results from the same volumetric analysis as done in [13] are shown in Figure 8 A). Overall, the same consistent decrease trend between HC and PwP as observed for the original GOUHFI for the putamen, hippocampus and amygdala was observed for GOUHFI 2.0-n2. It was only for putamen that all techniques presented a statistically significant difference between both HC and PwP sub-groups. For SynthSeg+, the median volumes measured were systematically larger than the rest. The p-values calculated are reported in Table 3. In addition, since TIV estimation is part of the GOUHFI 2.0 toolbox, an independent volumetry analysis with normalized volume values were performed without the use of SPM12. The same volumetric analysis, using TIV estimates from GOUHFI2.0-n2 and SynthSeg+, is shown in Figure 8 B), with corresponding p-values reported in Table 3. Albeit smaller than the normalized volumes reported using SPM12, the new normalized volumes resulted in the same statistical significance for the putamen. For both FastSurferVINN and the original GOUHFI, the TIV was not calculated independently due to differing definitions for the CSF label. An additional figure showing the volumetry results for the original GOUHFI, GOUHFI 2.0-n2 with TIV from SPM12 and GOUHFI 2.0-n2 with independent TIVs is provided in Appendix 6.2.
Figure 8: Box plots showing the normalized volumes computed by FastSurferVINN (left), GOUHFI 2.0-n2 (middle) and SynthSeg+ (right) for healthy controls (HC) and people with Parkinson’s disease (PwP) for the putamen, hippocampus and amygdala. In A), SPM12 was used for estimating the TIV for normalizing the volumes. In B), the same results are shown but GOUHFI 2.0-n2 and SynthSeg+ were used to estimate the TIV for each respectively (SPM12 was used for FastSurferVINN). For putamen, the three techniques had a statistically significant difference in volume between HC and PwP after Bonferroni correction for both A) and B).
@@ 4 Discussion
While testing the original GOUHFI [13], incidental observations of the limitations highlighted in Figure 4 prompted the idea of enhancing the diversity of the training dataset. One known limitation of the original training dataset was the absence of aged and/or non-healthy brain anatomies. The original GOUHFI exhibited limitations not observed for FastSurferVINN or SynthSeg+, likely due to the absence of non-healthy anatomies in its training dataset, which were included in the latter two. In order to address this issue, 32 aged and demented subjects were incorporated based on an ad hoc visual assessment of subjects from the OASIS1 dataset and only selecting subjects exhibiting substantially enlarged ventricles. Ultimately, segmentation errors like the ones observed in the original GOUHFI were completely resolved when qualitatively assessing GOUHFI 2.0’s output segmentations for the same test subjects, while also being absent in all other test subjects assessed.
Additionally, this limited capacity to delineate enlarged ventricles also highlighted a limitation of the generative model, which is its incapacity to generate synthetic images with anatomical features observed in non-healthy brains. Until further developments on the DR approach for brain segmentation are proposed, using an anatomically-varied training label maps is highly recommended. As outlined in [20], recent developments in diffusion and flow-based DL models show high potential for improving the image synthesis models, hopefully contributing to a more robust and varied synthesizing process.
@@ 4.2 SCAIFIELD: Comparing GOUHFI’s variants in healthy controls at 7T
Figure 5 presented the unexpected weakness of GOUHFI 2.0-n1 struggling with proper identification of posterior cerebellar cortex, exclusively observed for that model. Surprisingly, this suboptimal cerebellar cortex segmentation for GOUHFI 2.0-n1 did not seem to arise from typical inhomogeneities observed at UHF, since this was mostly observed in homogeneous N4-corrected pTx images (as in Figure 5) while not being observed in inhomogeneous 1 Tx 7T images. Ultimately, one possible explanation for this unexpected behavior for GOUHFI 2.0-n1 could be the randomness involvement in (1) the generation of synthetic images and (2) the initial seeding of the DL training process as shown in [1].
The superior identification of cerebellar WM branches by GOUHFI 2.0-n2 as shown for both subjects in Figure 5 makes it a tool of choice for researchers working with the cerebellum, outperforming FastSurferVINN and SynthSeg+. Improvements are also evident when comparing the original GOUHFI to the latest two variants of GOUHFI 2.0. As discussed in [13], even though SynthSeg+ also uses a DR approach, the difference in training resolutions (1.0 mm3 vs (0.7 mm)3) seems especially important for detecting small regions such as thin cerebellar WM branches.
@@ 4.3 SCAIFIELD-NPC: Assessment of segmentation methods in SCA subjects
Overall, no segmentation technique demonstrated a clear superiority over the others. It is important to mention that this section was focused on the cerebellum only. First, FastSurferVINN performed relatively well considering that these test cases were outside its optimized domain (i.e., (0.6 mm)3 7T images). It was the highest performing technique to properly delineate regions with substantial cortex atrophy. This improved delineation in atrophied cortical regions was most pronounced in the anterior lobe of the cerebellum where both SynthSeg+ and GOUHFI 2.0-n2 failed at accurately identifying the CSF-cortex border (cf. results for subject 1). For subjects 2 and 3 with less atrophy, this was less pronounced. This result can be surprising considering that both DR approaches were trained using perfectly aligned training image and label maps from the synthetic images produced by the generative model from the label maps. In other words, one could have expected a better identification of the CSF-GM border by SynthSeg+/GOUHFI 2.0 since the two regions have high contrast between each other. In the end, this circles back to the previous section discussing that a variation of anatomies inside the training corpus is still important, even when using synthetic training images. However, for subject 3, GOUHFI 2.0-n2 demonstrated a great capacity at accurately identifying the CSF-cortex border, except for the anterior lobe.
@@
On the other hand, FastSurferVINN considerably underestimated the inferior region of the posterior lobe of the cerebellar WM in two of the three test subjects while also systematically failing at properly delineating the inferior posterior cerebellar cortex, possibly due to the reduced signal due to inhomogeneities. Both issues can be problematic in volumetric studies where under-estimations of both WM and cortex volumes due to segmentation errors can bias the results. For SynthSeg+, the users should be aware of the opposite trend: over-estimation of WM and cortex volumes. As discussed in the last paragraph, SynthSeg+ wrongly segmented a substantial amount of cortex and CSF as cerebellum WM. For the cortex, SynthSeg+ performed the worst at identifying small atrophy regions, resulting in the cortex label to be segmented as a oversimplified continuous shape instead of following the ”branchy” nature of the atrophied cerebellum cortex for SCA subjects. For GOUHFI2.0-n2, the marginal WM overestimation and improved cortex–CSF border identification, compared to SynthSeg+, indicate that it is a better alternative for high-resolution UHF images.
Regarding the mislabeling by GOUHFI 2.0 of the cerebellar cortex with cerebral WM (cf. Subject 2 in Figure 6), a pathological hallmark of SCAs is the degeneration of the so-called Purkinje-cell layer, leading to a gliotic aspect of affected cerebellar areas and thereby a thinning of the cerebellar cortex. The anterior lobe of the cerebellum is known to be particularly affected in the most frequent types of SCA in central and northern Europe presented here, sometimes even before the onset of motor symptoms. These changes are known to be visible in structural MRI and are used diagnostically by clinicians. It is therefore noteworthy that we observed particular difficulties of GOUHFI 2.0 to segment the cerebellar cortex of the anterior lobe. This might indicate that the tool’s inability to identify cortical areas of the cerebellum here, points to the tool accidentally picking up an important pathophysiological feature of the disease[11, 44, 36].
@@ 4.4 TIV estimation
One interesting result from the TIV estimations was that the only cases where the GOUHFI 2.0 variants under-estimated the TIV, compared to SPM12, are the two HCP contrasts. In contrast, SynthSeg+ exhibited a systematic overestimation of the TIV for the HCP dataset and, more broadly, across all ten test datasets. One difference compared to all other test datasets was that the brain extraction step was not performed as part of this work, since it was already performed as part of the HCP preprocessed data. In both cases, the brain extraction process was very accurate, with no inclusion of extra tissue. For STRAT-PARK, only brain-extracted images were provided, with Nighres [21] used for the brain extraction step. For most subjects in that dataset, a substantial amount of extra-cerebral tissue was still present in the brain-extracted images as shown in Figure 4. For the remaining datasets, the brain extraction step was performed with the ANTsPyNet brain extraction tool implemented as part of the GOUHFI toolbox. The results were visually assessed, and cases with minor extra-cerebral tissue were still retained (negligible compared to START-PARK), as this had no impact on brain segmentation aside from the CSF label. Indeed, for GOUHFI2.0-n2, visual inspection revealed that the CSF label often incorporated extra-cerebral tissue when brain extraction quality was low. On the other hand, GOUHFI 2.0-n1 was not as affected by extra-cerebral tissue, resulting in less TIV overestimation. While the GOUHFI 2.0-n1 variant resulted in higher-quality TIV estimates than GOUHFI 2.0-n2, the latter was still selected as the superior variant due to the reasons stated in section 4.2.
@@
While FastSurferVINN was not considered the reference for earlier brain segmentation quantitative comparison in this study, it was selected as the reference for the volumetry analysis. Based on extensive visual inspection of the segmentation results, FastSurferVINN was deemed as the technique with the most accurate delineation of the hippocampus-ventricle laterally. However, in subjects with strong signal inhomogeneities, medial portions of both the hippocampus and amygdala may be excluded, causing lower estimated volumes and possibly explains the lower estimates in Figure 8 A) compared to GOUHFI 2.0-n2 and SynthSeg+.
@@ 4.7 Limitations
The main limitation of the GOUHFI 2.0 toolbox is the brain extraction requirement. As already discussed in the original publication of GOUHFI, most segmentation techniques do not share this requirement and is partly explained by lack of robust whole-head segmentation techniques for UHF images [13]. Future advances in whole-head segmentation algorithms for UHF images with low signal outside the head, or the use of generative diffusion models to generate synthetic head labels, could help address this issue.